OpenStudy (babycease):

he table shows the outputs y for different inputs x: Input (x) 5 6 7 8 Output (y) 2 5 8 11 Part A: Do the data in this table represent a function? Justify your answer. (3 points) Part B: Compare the data in the table with the relation f(x) = 2x + 13. Which relation has a greater value when x = 7? (2 points) Part C: Using the relation in Part B, what is the value of x if f(x) = 75? (5 points) Help Please will give medal

Mehek (mehek14):

Use the slope formula for part A \(\dfrac{6-5}{5-2}\)

OpenStudy (babycease):

you get 1/2

Mehek (mehek14):

Not 1/2

Mehek (mehek14):

5-2 = ?

OpenStudy (babycease):

so sorry three

OpenStudy (babycease):

1/3

Mehek (mehek14):

My bad Flip it so it would be 3

Mehek (mehek14):

Should be \(\dfrac{5-2}{6-5}\)

Mehek (mehek14):

Next 2 ordered pairs \(\dfrac{11-8}{8-7}=?\)

OpenStudy (babycease):

3/1

Mehek (mehek14):

so both are 3 that's a function

Mehek (mehek14):

So for Part A put this: Yes it is a function because the rate of change is 3

Mehek (mehek14):

For part B, put 7 instead of x f(7) = 2 * 7 + 13 = ?

OpenStudy (babycease):

okay thank you im understanding better

OpenStudy (babycease):

the answer would be 27

Mehek (mehek14):

Ok so when x = 7, y = 27 in the other function, when x = 7, y = 8

Mehek (mehek14):

Which one is greater when x = 7?

OpenStudy (babycease):

um i suppose when y is 27

Mehek (mehek14):

Correct :)

Mehek (mehek14):

So for Part B put: the relation f(x) = 2x + 13 is greater when x = 7

OpenStudy (babycease):

okay last one

Mehek (mehek14):

For that, we put 75 instead of f(x) and solve for x 75 = 2x + 13

Mehek (mehek14):

Any idea how to solve?

OpenStudy (babycease):

um get x by itself

Mehek (mehek14):

Yes How would we do that?

OpenStudy (babycease):

Not sure tbh i think subtract 75 and 13

Mehek (mehek14):

good 75 - 13 = ?

OpenStudy (babycease):

is 62

Mehek (mehek14):

Yep now divide by 2

OpenStudy (babycease):

31

Mehek (mehek14):

so the value of x is 31

OpenStudy (babycease):

so how would it look

Mehek (mehek14):

For part c, just put When the value of x is 31

OpenStudy (babycease):

ok